# Information scrambling and entanglement dynamics in Floquet Time Crystals

### v1

@@ II The Model
$\hat{U}_{F}=\exp(-i\hat{H}_{0}T)\,\hat{U}_{K},\quad\hat{U}_{K}\equiv\exp\left(i\phi\sum_{i=1}^{N}\hat{\sigma}_{i}^{x}/2\right),$ (2)
@@ III FTC dynamics
In this section, we show our results for the dynamics of magnetization, OTOCs, and entanglement entropy in the FTC phase. Their behaviour follow different growth/decay rates at corresponding characteristic time scales, which depend on the coupling parameters and system size, as we discuss in detail. In all our simulations we consider initial product states $\left[\cos\theta\lvert\uparrow\rangle+\sin\theta\lvert\downarrow\rangle\right]^{\otimes N}$, with $\theta=\pi/8$, open boundary conditions and average our results over $10^{3}$ disorder simulations.
@@ III.1 Magnetization dynamics
takes a constant value. Most importantly, for large system sizes $N$ and disordered interactions $J_{i}$, the period doubling dynamics is stable under perturbations of the Floquet operator such as a non-zero transverse field $h^{x}$ or imperfect spin flip $\phi=\pi+\epsilon$ [2]. Therefore, the system breaks the time-translation symmetry stabilizing a FTC phase.
@@
$\bullet\ \,\mathcal{T}_{1}<t<\mathcal{T}_{2}$: a stable period doubling dynamics till time $\mathcal{T}_{2}$, with all the spins now in a coherent dynamics due to kicking and MBL interactions. Due to MBL, the time scale for period doubling dynamics increases exponentially with the system size, $\mathcal{T}_{2}\sim e^{\beta N},\beta>0$, therefore persisting indefinitely in the macroscopic limit, $N\rightarrow\infty$.
$\bullet\ \,\mathcal{T}_{2}<t<\mathcal{T}_{3}$: for longer times, the spins now start to decohere among each other and the magnetization $Z(t)$ slowly decays. We find a logarithmic decay leading to a characteristic time scale exponentially long with the system size, $\mathcal{T}_{3}=e^{\beta^{\prime}N}$. In particular, for the chosen parameters, a time exponent even larger than the period doubling regime, with $\beta^{\prime}\approx 0.95>\beta\approx 0.43$.
@@ III.2 Out-of-time-ordered correlator
where $\ell=1,...,N$, and $\alpha,\beta\in\{x,y,z\}$. We focus on $C_{xx}(\ell,t),C_{zz}(\ell,t)$ as these are representative of the main behaviors of the OTOC in the system.
@@ III.3 Entanglement dynamics
$S=-\frac{\text{Tr}\left[\hat{\rho}_{A}\log(\hat{\rho}_{A})\right]}{L_{A}}$ (8)
where $L_{A}=N/2$ is size of subset $A$. In a thermal system, the entanglement entropy of an initial low entangled state typically increases linearly over time until it reaches the value associated with the system’s energy density, where the entanglement entropy obeys the thermodynamic volume law. Conversely, in a localized system, one may expect that since the information remains confined near the subsystem, the entanglement entropy should saturate at a constant value proportional to the localization length rather than the subsystem size. In fact, in AL system, the entanglement entropy saturates to a constant independent of the system size. In generic MBL systems, however, it has been observed [87, 88, 89, 79] that the entanglement entropy grows slowly, following a $\log t$ dependence in time, and saturates to a volume law at exponentially long times.

### v2

@@ II The Model
$\hat{U}_{F}=\exp(-i\hat{H}_{0}T)\,\hat{U}_{K},\quad\hat{U}_{K}\equiv\exp\left(i\phi\sum_{i=1}^{L}\hat{\sigma}_{i}^{x}/2\right),$ (2)
@@ III FTC dynamics
In this section, we show our results for the dynamics of magnetization, OTOCs, and entanglement entropy in the FTC phase. Their behaviour follow different growth/decay rates at corresponding characteristic time scales, which depend on the coupling parameters and system size, as we discuss in detail. In all our simulations we consider initial product states $\left[\cos\theta\lvert\uparrow\rangle+\sin\theta\lvert\downarrow\rangle\right]^{\otimes L}$, with $\theta=\pi/8$, open boundary conditions and average our results over $10^{3}$ disorder simulations. However, our results are independent of the choice of initial state.
@@ III.1 Magnetization dynamics
takes a constant value. Most importantly, for large system sizes $L$ and disordered interactions $J_{i}$, the period doubling dynamics is stable under perturbations of the Floquet operator such as a non-zero transverse field $h^{x}$ or imperfect spin flip $\phi=\pi+\epsilon$ [2]. Therefore, the system breaks the time-translation symmetry stabilizing a FTC phase.
@@
$\bullet\ \,\mathcal{T}_{1}<t<\mathcal{T}_{2}$: a stable period doubling dynamics till time $\mathcal{T}_{2}$, with all the spins now in a coherent dynamics due to kicking and MBL interactions. Due to MBL, the time scale for period doubling dynamics increases exponentially with the system size, $\mathcal{T}_{2}\sim e^{\beta L},\beta>0$, therefore persisting indefinitely in the macroscopic limit, $L\rightarrow\infty$.
$\bullet\ \,\mathcal{T}_{2}<t<\mathcal{T}_{3}$: for longer times, the spins now start to decohere among each other and the magnetization $Z(t)$ slowly decays. We find a logarithmic decay leading to a characteristic time scale exponentially long with the system size, $\mathcal{T}_{3}=e^{\beta^{\prime}L}$. In particular, for the chosen parameters, a time exponent even larger than the period doubling regime, with $\beta^{\prime}\approx 0.95>\beta\approx 0.43$.
@@ III.2 Out-of-time-ordered correlator
where $\ell=1,...,L$, and $\alpha,\beta\in\{x,y,z\}$. We focus on $C_{xx}(\ell,t),C_{zz}(\ell,t)$ as these are representative of the main behaviors of the OTOC in the system.
@@ III.3 Entanglement dynamics
$S=-\frac{\text{Tr}\left[\hat{\rho}_{A}\ln(\hat{\rho}_{A})\right]}{L_{A}}$ (8)
where $L_{A}=L/2$ is size of subset $A$. In a thermal system, the entanglement entropy of an initial low entangled state typically increases linearly over time until it reaches the value associated with the system’s energy density, where the entanglement entropy obeys the thermodynamic volume law. Conversely, in a localized system, one may expect that since the information remains confined near the subsystem, the entanglement entropy should saturate at a constant value proportional to the localization length rather than the subsystem size. In fact, in AL system, the entanglement entropy saturates to a constant independent of the system size. In generic MBL systems, however, it has been observed [87, 88, 89, 79] that the entanglement entropy grows slowly, following a $\log t$ dependence in time, and saturates to a volume law at exponentially long times.
@@ Appendix A Exact OTOC evolution
Figure 5: The dynamics associated with melted FTC (ergodic phase) for system-size $L=8$ and parameter values $J=1=h^{z}$, $h^{x}=0.5$ and the kicking phase $\phi=\pi/2+0.4$. (Left) The magnetization $Z(t)$, (center) the OTOCs along the $z$-direction ($C_{zz}(l,t)$) and (right) $x$-direction ($C_{xx}(l,t)$), for varying spin distances $l$. The inset shows the early time growth.
@@
Figure 6: (a) The entanglement entropy $S(t)$ in melted FTC phase for the varying system-sizes $L$, and parameter values $J=1=h^{z}$, $h^{x}=0.5$ and the kicking phase $\phi=\pi/2+0.4$. (b) The saturation time plotted as function of system-size $L$, where the dash-dotted lines shows the fitted curve $\sim L^{\alpha}$ ($\alpha\approx 0.87$). (c) The saturation value $S(\infty)$ as a function of $L$ shows that the finite-size effect decays with system-size. The dash-dotted lines show the fitted curve $S^{(N\rightarrow\infty)}(\infty)+a/L$ with $S^{(N\rightarrow\infty)}(\infty)\approx 0.65$.
@@
Appendix B Ergodic phase
In this section, we examine the dynamics in the melted FTC or ergodic phase by introducing a phase kick $\phi$ far away from $\pi$. For our numerical simulations, we set the coupling parameters to $J=h^{z}=1$, $h^{x}=0.5$, and the $\phi=\pi/2+0.4$ for a system size of $L=8$. Under these parameters, the mean level spacing ratio $r$ of the quasi-Hamiltonian $i\ln\hat{U}_{F}$ (associated with the unitary evolution in Eq. (2)) is computed to be $\approx 0.53$ indicating a Wigner-Dyson distribution, typical of ergodic systems. As shown in Fig. 5 (left), the magnetization rapidly decays to zero due to thermalization. Unlike the FTC phase, the melted phase lacks a preferred protected direction, leading to identical scrambling behaviour along all directions. This is evident in Fig. 5 (center and right), where the OTOCs $C_{zz}(l,t)$ and $C_{xx}(l,t)$ exhibit identical behavior. Similar to thermal systems, we observe a rapid early growth followed by saturation. In early growth regime ($C_{\alpha\beta}(l,t)<<1$), the behavior obeys the previously proposed conjecture [72]
$C_{\alpha\beta}(l,t)\sim\exp(-\lambda(l-v_{B}t)^{1+p}/t^{p})\,.$ (16)
Here, $v_{B}$ is the butterfly velocity, and $p$ is the wave front broadening coefficient. In Fig. 6a, we present the behavior of entanglement entropy as a function of system size $L$, keeping the coupling parameters fixed. The entanglement entropy initially grows before reaching saturation. Unlike the FTC phase, where the saturation time exhibits exponential scaling, in the melted FTC, it scales polynomialy with system size $t_{\text{sat}}\sim L^{\alpha}$(see Fig. 6b). The saturation time is determined by first computing the saturation value $S(\infty)$ and then identifying the earliest time $t_{\text{sat}}$ at which $|S(t_{\text{sat}})-S(\infty)|\leq\epsilon$ where $\epsilon$ is a tolerance parameter set to $10^{-3}$ in our numerics. Due to finite-size effects, the saturation value $S(\infty)$ is not constant for the considered system sizes, as expected in a thermal system. We perform a finite-size scaling analysis of these saturation values to show that the correction decays to zero as $L\rightarrow\infty$ (see Fig. 6c).